# Fusion rule in conformal field theories and topological orders: A unified view of correspondence and (fractional) supersymmetry and their relation to topological holography

### v5

@@ III.1 Vanishing of fusion rule and (fractional) supersymmetry
In the Majorana-Ising CFT, by the chiral (or antichiral) semion $\{e,m\}$ representation, one can identify the order and disorder operator in $\mathbf{F}^{(0)}$ as $\sigma_{\text{Bulk}}=(e\overline{e}+m\overline{m})/\sqrt{2},\mu_{\text{Bulk}}=(m\overline{e}+e\overline{m})/\sqrt{2}$[23, 27]. Hence, one can reproduce the bulk fusion rule of the original bulk CFT from the combination of the chiral fields, $\{\phi\}=\{I,\psi,e,m\}=\{\mathbf{I},\mathbf{f},\mathbf{e},\mathbf{m}\}=\mathbf{S}$, and their antichiral counterparts.
@@
This degenerate fusion rule implies that the two representations live in different Hilbert spaces or lattice models.
@@ IV Relationship between our formalism and existing category theories
In this section, we clarify the precise relationship between our algebraic formalism and existing category theories. Our construction can be summarized as a construction of a $Z_{N}$ graded fusion algebra (or $Z_{N}$ graded CCFT) from $Z_{N}$ symmetric SFC or $Z_{N}$ symmetric bosonic full CFT. We itemize the construction of the $Z_{N}$ extended CCFT in the previous sections using the terminology in category theories as follows.
  • Choosing spherical fusion category $\mathbf{B}$ with $Z_{N}$ action (or $Z_{N}$ simple current). This corresponds to a bosonic modular invariant and MTC[114, 116] with a connection to the Verlinde formula[5, 6].
@@
We summarize the relationship between our formalism and the related category theories in FIG.1. The resultant $Z_{N}$ graded chiral algebra $\{\phi\}$ can be considered as $Z_{N}$ generalization of the superfusion-category [181, 182, 183, 184, 185], or the “fractional superfusion category", but further investigations are necessary. (Interestingly, the introduction of $\phi_{a,p}$ results in the emergence of entangled paired particles when considering the chiral correlation functions or corresponding wavefunctions[186, 173, 174, 175, 27].)
@@
In this work, we have studied the structure of fusion rules in CFTs with a modern view and clarified a unified relationship between bulk CFT, CCFT, and TO. Especially, the relationship between the bulk and chiral Majorana-Ising CFT and the double semion algebra in the $2+1$-dimensional TO has been established directly by studying the chiral and bulk semionization of the model. Moreover, we have found a new structure, the vanishing fusion rule, and this can explain the duality or (fractional) supersymmetry appearing in contemporary physics in a unified and concise way. Further investigation into this structure on lattice models is an interesting future problem. We expect that the same procedure can apply to CFTs and the corresponding TOs in general space-time dimensions and modify the recent analysis on topological holography (FIG. 2). By combining the Moore-Seiberg data[56] and symmetry-based analysis of RG domain wall[195] in [73], our method will enable one to unify the ideas in TOs and their gauging structure [^6] After the submission of the first version of the present work, we noticed interesting applications of CCFT[27] to the TOs of mixed quantum states[173, 174, 175, 176]. However, it should be noted that their models are still in the scope of the CFT/TQFT combined with the Witt equivalence. This implies the fundamental importance of studying anyons and generalized symmetry in manipulating general mixed quantum states. Related quantum information theoretical aspects which might be useful for this problem can be seen in [211], for example.. We also note the existence of the exceptional models outside of the Moore-Seiberg data in [197, 60] for further studies.
@@ Appendix A Basis transformations and symmetry in lattice models
$(\sigma\overline{\sigma})\times(\sigma\overline{\sigma})=I+\psi+\overline{\psi}+\epsilon.$ (61)

### v7

@@ III.1 Vanishing of fusion rule and (fractional) supersymmetry
In the Majorana-Ising CFT, by the chiral (or antichiral) semion $\{e,m\}$ representation, one can identify the order and disorder operator in $\mathbf{F}^{(0)}$ as $\sigma_{\text{Bulk}}=(e\overline{e}+m\overline{m})/\sqrt{2},\mu_{\text{Bulk}}=(m\overline{e}+e\overline{m})/\sqrt{2}$[23, 27]. Hence, one can reproduce the bulk fusion rule of the original bulk CFT from the combination of the chiral fields, $\{\phi\}=\{I,\psi,e,m\}=\{\mathbf{I},\mathbf{f},\mathbf{e},\mathbf{m}\}=\mathbf{S}$, and their antichiral counterparts. We also note that the expression $\sigma_{\text{Bulk}}=\sigma\boxtimes\overline{\sigma}$ can lead to the naive application of the combinations of the fusion and Deligne products resulting in the wrong expression $\psi\times\sigma_{\text{Bulk}}=\psi\times(\sigma\boxtimes\overline{\sigma})=\sigma_{\text{Bulk}}$. It should be stressed that the application of the chiral fermion $\psi$ to the bulk order operator $\sigma_{\text{Bulk}}$ should produce the bulk disorder operator $\mu_{\text{Bulk}}$, and this fact has been established recursively in the studies of both QFT and statistical mechanical models in common. Our expressions reproduce the established result, counting of order and disorder operators, and other existing fusion rules of the Majorana theory in a concrete and rigorous way.
@@
This degenerate fusion rule implies that the two representations live in different Hilbert spaces or lattice models. To our knowledge, this degenerate fusion rule has not been studied in literature.
@@ IV Relationship between our formalism and existing category theories
In this section, we clarify the precise relationship between our algebraic formalism and existing category theories. Our construction can be summarized as a construction of a $Z_{N}$ graded fusion algebra (or $Z_{N}$ graded CCFT) $\mathbf{S}$ from $Z_{N}$ symmetric SFC or $Z_{N}$ symmetric bosonic full CFT $\mathbf{B}$. $\mathbf{S}$ provides a new type of of $Z_{N}$ graded algebras corresponding to $Z_{N}$ graded SymTFTs, and they correspond to the modular partition function Eq. (13) which is absent (or at least less common) in literature. We itemize the construction of the $Z_{N}$ extended CCFT in the previous sections using the terminology in category theories as follows.
  • Choosing spherical fusion category $\mathbf{B}$ with $Z_{N}$ action (or $Z_{N}$ simple current). This corresponds to a bosonic modular invariant and MTC $\mathbf{M}$ [126, 128] with a connection to the Verlinde formula[5, 6].
@@
We summarize the relationship between our formalism and the related category theories in FIG.1. The resultant $Z_{N}$ graded chiral algebra $\{\phi\}$ can be considered as $Z_{N}$ generalization of the superfusion-category [190, 191, 192, 193, 194], or the “fractional superfusion category", but further investigations are necessary. (Interestingly, the introduction of $\phi_{a,p}$ results in the emergence of entangled paired particles when considering the chiral correlation functions or corresponding wavefunctions[195, 183, 184, 185, 27].) We also remind again that the procedure of taking a subalgebra corresponds to the application of the massive RG flow, and this reduces a bulk CFT to (smeared) BCFT[65, 66, 67], which corresponds to CCFT under the open-closed duality[5, 6]. The bulk semionization provides a concrete and concise algebraic representation of this phenomenon, which contrasts with existing analytical methods.
@@
In this work, we have studied the structure of fusion rules in CFTs with a modern view and clarified a unified relationship between bulk CFT, CCFT, and TO. Construction of the $Z_{N}$ graded SymTFTs, $\mathbf{S}$, and the analysis of the corresponding partition function (13) are the main results. Especially, the relationship between the bulk and chiral Majorana-Ising CFT and the double semion algebra in the $2+1$-dimensional TO has been established directly by studying the chiral and bulk semionization of the model. Moreover, we have found a new structure, the vanishing fusion rule, and this can explain the duality or (fractional) supersymmetry appearing in contemporary physics in a unified and concise way. Further investigation into this structure on lattice models is an interesting future problem. We expect that the same procedure can apply to CFTs and the corresponding TOs in general space-time dimensions and modify the recent analysis on topological holography (FIG. 2). By combining the Moore-Seiberg data[68] and symmetry-based analysis of RG domain wall[204] in [85], our method will enable one to unify the ideas in TOs and their gauging structure [^7] After the submission of the first version of the present work, we noticed interesting applications of CCFT[27] to the TOs of mixed quantum states[183, 184, 185, 186]. However, it should be noted that their models are still in the scope of the CFT/TQFT combined with the Witt equivalence. This implies the fundamental importance of studying anyons and generalized symmetry in manipulating general mixed quantum states. Related quantum information theoretical aspects which might be useful for this problem can be seen in [221], for example.. We also note the existence of the exceptional models outside of the Moore-Seiberg data in [206, 72] for further studies.
@@ Appendix A Basis transformations and symmetry in lattice models
$(\sigma\otimes\overline{\sigma})\times(\sigma\otimes\overline{\sigma})=I+\psi+\overline{\psi}+\epsilon.$ (61)
@@ Appendix B Gauging and condensation in coupled models: emergence of noninvertible or nonabelian objects
Appendix C Bulk semionization for $SU(3)_{3}$ WZW model
In this section, we apply our method to $SU(3)_{3}$ WZW model with integer spin $Z_{3}$ simple currents. First, we note the data of the chiral fusion rule or MTCs as $\mathbf{M}=\{\phi_{o,p},\phi_{j,p},\phi_{j^{\dagger},p},\theta_{a}\}$, where the conformal dimensions are summarized as follows:
$Q=0$ $:h_{o,0}=0,h_{o,1}=h_{o,2}=1,h_{a}=1/2,$ (66)
$Q=1/3$ $:h_{j,0}=2/9,:h_{j,1}=5/9,h_{j,2}=8/9,$ (67)
$Q=2/3$ $:h_{j^{\dagger},0}=2/9,:h_{j^{\dagger},2}=5/9,h_{j^{\dagger},1}=8/9,$ (68)
where $Q$ is the $Z_{3}$ charge specifying the sectors. In this model, the $Z_{3}$ simple currents corresponds to $\phi_{o,1}$ and $\phi_{o,2}$, and only the field $\theta_{a}$ is $Z_{3}$ invariant. Corresponding to this structure, one can identify the lower index $p=0,1,2$ as $Z_{N}$ parity. Hence, we focus on the $Z_{3}$ invariant object $\theta_{a}$ and its splitting under the extension and bulk semionization. Some part of the fusion rules with charge cancellation can be seen in [220], and the more general fusion rules are in the scope of the Verlinde formula[5]. Hence, with some patient calculations, one can obtain whole fusion rules. To simplify the discussion, we mainly focus on the fusion rule $\theta_{a}\times\theta_{a}=\sum_{p}\phi_{o,p}+2\theta_{a}$.(In other words, we demonstrate that one can calculate some parts of the extended fusion rules from the given algebraic data without information on the all algebraic data.)
Under the Moore-Seiberg data[68, 69], one can obtain the nonchiral fusion ring through the ring isomorphism, $\mathbf{B}=\{\Phi_{o,p,-p},\Phi_{j,p,-p},\Phi_{j^{\dagger},p,-p},\Phi_{a,0}\}=\{\phi_{o,p},\phi_{j,p},\phi_{j^{\dagger},p},\theta_{a}\}=\mathbf{M}$. By applying the $Z_{3}$ extension, the theory can be represented as $\mathbf{F}=Z_{N}\otimes\mathbf{B}=\{\Phi_{o,p,\overline{p}},\Phi_{j,p,\overline{p}},\Phi_{j^{\dagger},p,\overline{p}},\Phi_{a,p}\}$. Hence, the relation $\theta_{a}\times\theta_{a}=\sum_{p}\phi_{o,p}+2\theta_{a}$ is modified as follows,
$\Phi_{a,p}\times\Phi_{a,p^{\prime}}=\sum_{p^{\prime\prime}}\Phi_{o,p^{\prime\prime}+p+p^{\prime},-p^{\prime\prime}}+2\Phi_{a,p+p^{\prime}}$ (69)
By the bulk semionization, one can introduce the operators in the $Z_{3}$-graded SymTFTs $\mathbf{S}$ as $\Psi_{a,p}=\Phi_{a,p}/\sqrt{3}$, $\Psi_{o,p}=\left(\sum_{p^{\prime}}\Phi_{o,p^{\prime}+p,-p^{\prime}}\right)/3$. Hence, after the bulk semionization $\mathbf{S}\subset\mathbf{F}$, just by deviding the left and right handside of Eq.(69) by $3$, one can obtain the resultant fusion rule in $\mathbf{S}$ as
$\Psi_{a,p}\times\Psi_{a,p^{\prime}}=\Psi_{o,p+p^{\prime}}+\frac{2}{\sqrt{3}}\Psi_{a,p+p^{\prime}}.$ (70)
Interestingly, the coefficient $2/\sqrt{3}$ inevitably appears. We stress that, if one does not permit the appearance of such noninteger coefficients, the algebraic descriptions of the sandwich construction fail. Moreover, in [161], the author demonstrated that even in the $SU(2)_{4}$ WZW model, the simplest model with the $Z_{2}$ bosonic simple current, this kind of unusual fusion coefficient appears. The precise relationship between the unusual fusion coefficients and the boson-fermion statistics may be an interesting problem.
By applying the same method to other fusion rules involving the multiplication of $\theta_{a}$, one can obtain the corresponding fusion rules of the $Z_{3}$-graded SymTFTs. For the other fusion rules only involving $\phi$ of the MTC, the fusion rules are unchanged by the bulk semionization.